# Cost-Effective Text Clustering with Large Language Models

### v1

@@ Abstract.
Text clustering aims to automatically partition a collection of text documents into distinct clusters based on linguistic features. In the literature, this task is usually framed as metric clustering based on text embeddings from pre-trained encoders or a graph clustering problem upon pairwise similarities from an oracle, e.g., a large ML model. Recently, large language models (LLMs) bring significant advancement in this field by offering contextualized text embeddings and highly accurate similarity scores, but meanwhile, present grand challenges to cope with substantial computational and/or financial overhead caused by numerous API-based queries or inference calls to the models.
In response, this paper proposes TECL, a cost-effective framework that taps into the feedback from LLMs for accurate text clustering within a limited budget of queries to LLMs. Under the hood, TECL adopts our EdgeLLM or TriangleLLM to construct must-link/cannot-link constraints for text pairs, and further leverages such constraints as supervision signals input to our weighted constrained clustering approach to generate clusters. Particularly, EdgeLLM (resp. TriangleLLM) enables the identification of informative text pairs (resp. triplets) for querying LLMs via well-thought-out greedy algorithms and accurate extraction of pairwise constraints through carefully-crafted prompting techniques. Our experiments on multiple benchmark datasets exhibit that TECL consistently and considerably outperforms existing solutions in unsupervised text clustering under the same query cost for LLMs.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Information systems Clusteringccs: Information systems Language models
@@ 1. Introduction
With the advent of LLMs, another way for text clustering is to obtain similarity scores of text pairs from LLMs to define a graph clustering problem (Silwal et al., 2023). This methodology enables us to model the correlations between text instances with highly accurate semantic similarities/dissimilarities from LLMs, and hence, producing high-quality clustering results (Silwal et al., 2023; Zhang et al., 2023a). Despite the superb performance achieved, it requires pair-wise similarities between all text instances, leading to $O(n^{2})$ inference calls to the models, where $n$ is the total number of texts in the input text corpus, which is infeasible for clustering large-scale text collections in practice. Very recently, there are a number of works (Zhang et al., 2023a; Wang et al., 2023; Yang et al., 2024; Huang and He, 2024) exploring other ways of leveraging the feedback from instruction-tuned LLMs to guide text clustering. However, all of them still demand substantial queries to LLMs (typically with a token count of over 10×10$\times$10 × the corpus size) and/or rely on skeleton support from fine-tuning embedders such as BERT, Instructor, and E5, which is time-consuming.
@@
Motivated by this, this paper proposes TECL (short for cost-effective TExt Clustering with LLMs), which aims to maximize the text clustering quality within a fixed budget (i.e., for token consumption) for querying LLMs. To be more specific, TECL is inspired by the constrained clustering paradigm (a.k.a. semi-supervised clustering) (Dinler and Tural, 2016), which additionally takes as input a number of constraints for pairs of data points, e.g., must-links or cannot-links, from human experts. As such, clusters can thus be obtained based on these constraints as supervision signals and cheap text embeddings from small PLMs, e.g., BERT or Sentence-BERT (Reimers, 2019). Given the human-level performance of LLMs in text comprehension, a simple and straightforward approach is to employ LLMs as the domain expert to create pair-wise pseudo constraints. However, this scheme poses two pressing challenges. First, it still remains unclear which text instances should be selected for querying and how to elicit high-caliber responses from LLMs. Secondly, the answers from LLMs might be noisy or random, engendering unreliable constraints that mislead the subsequent clustering.
@@
To tackle the above problems, TECL includes two approaches EdgeLLM and TriangleLLM that both construct must-link and cannot-link constraints with LLMs in a cost-effective manner. EdgeLLM harnesses the prominent spanning edge centrality (Teixeira et al., 2013; Chandra et al., 1989) in graph theory to quantify the informativeness of text pairs (referred to as edges) and a well-thought-out greedy algorithm to achieve a linear-complexity identification of crucial edges. For each selected edge $(t_{a},t_{b})$, EdgeLLM elicits the relation, e.g., must-link (must be in the same cluster) or cannot-link (cannot belong to the same cluster), of $t_{a}$ and $t_{b}$ from LLMs with carefully-crafted prompts containing the task context and text contents. Based thereon, we further develop TriangleLLM that greedily selects crucial text triplets (referred to as triangles) for querying, which draws inspiration from the fact that humans represent an instance through comparing with others (Nosofsky, 2011). On top of that, TECL refines classic constrained clustering approaches, i.e., constrained spectral clustering (Wang et al., 2014) and constrained K-Means (Wang et al., 2014), as WCSC and WCKMeans, with our carefully designed penalty weights for pairwise constraints based on the pointwise mutual information (Church and Hanks, 1990) calculated from the text embeddings, so as to eliminate the side effects of misinformation introduced by LLMs.
@@
Our extensive experiments comparing TECL against multiple strong baselines under the same query budget for LLMs on five benchmark text datasets demonstrate the consistent and superior performance of TECL in unsupervised text clustering tasks. In particular, on the well-known Tweet dataset, TECL is able to attain an accuracy of $73.1\%$ with a remarkable improvement of $7.29\%$ over the state-of-the-art at a monetary cost of \$0.0185 for calling GPT-4.0-Mini.
@@ 2.1. Text Clustering
Early methods for text clustering primarily rely on vector space models, e.g., bag-of-words and TF-IDF (Schütze et al., 2008), where documents are represented as high-dimensional sparse vectors and clustered using k-Means, DBSCAN, or other metric clustering algorithms. By transforming words into low-dimensional dense vectors based on their relationships in corpora, Word2Vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) efficiently produce text embeddings encoding semantic and syntactic similarities, but still fail to capture contextual variations and polysemy. Transformer-based language models, e.g., BERT (Devlin, 2018) and its successors (Liu, 2019; Reimers, 2019; Su et al., 2023), mitigate the above issues by generating contextual embeddings via pre-training on a large corpus, which have set new benchmarks in text clustering.
Although the aforementioned embedding-based methods are simple and efficient, they fail to capture deeper semantic nuances in complex domains (Sun et al., 2019; Saha, 2023; Ravi and Kulkarni, 2023; Viswanathan et al., 2023; Muennighoff et al., 2023). A series of deep clustering approaches  (Zhou et al., 2022) are proposed to learn representations and cluster assignments jointly. They can be categorized into multiple groups, such as multi-stage (Tao et al., 2021; Huang et al., 2014), iterative (Yang et al., 2016; Caron et al., 2018; Van Gansbeke et al., 2020; Niu et al., 2022; Chang et al., 2017; Niu et al., 2020), generative (Dilokthanakul et al., 2016), and simultaneous methods (Xie et al., 2016; Zhang et al., 2021a; Hadifar et al., 2019), etc. While being effective, these approaches typically demand training on large datasets and careful tuning.
Recently, inspired by the powerful natural language capabilities of LLMs, several efforts have been made towards harnessing LLMs for enhanced text clustering. In Ref. (Keraghel et al., 2024; Petukhova et al., 2025), researchers have validated the superiority of LLMs over other text encoders when using their output text embeddings for clustering. Wang et al. (2023) propose to classify samples with explanations for clusters generated from LLMs, while De Raedt et al. (2023) collect descriptive labels using prototypical utterances to support in-context learning. ClusterLLM (Zhang et al., 2023a) achieves impressive clustering quality with sentence relations predicted by LLMs as the guide and finetuned embedders. Based on a subset of ground truth labels or expert feedback, Viswanathan et al. (2023) expand keyphrases and refine low-confidence cluster assignments to facilitate semi-supervised clustering, whose process pipeline is most similar to ours if reserving the LLM oracle and dropping other components. LLMEdgeRefine (Feng et al., 2024) creates superpoints to mitigate outliers and iteratively refines clusters using LLMs for improved semantic coherence. Yang et al. (2024) fine-tune embedding models based on the weakly-supervised contrastive learning loss and feedback from LLMs, followed by using LLMs to refine cluster assignments of low-confidence samples. Huang and He (2024) propose to transform the text clustering into a classification task by prompting LLMs to generate potential labels and assign the most appropriate label to each sample. Despite encouraging results reported, prior works strongly rely on extensive LLM queries, undergo costly fine-tuning, and/or require additional post-corrections, rendering them rather inefficient and impractical. Unlike previous works, our method achieves superior performance under identical conditions with a sufficiently low number of queries and budgets without fine-tuning or post-correction.
@@ 2.2. Constrained Clustering
To improve clustering performance, constrained clustering adopts side information such as pairwise must-link (ML) or cannot-link (CL) constraints. Classical approaches incorporate pairwise constraints into traditional clustering algorithms by refining their optimization frameworks. For example, constraint-based variants of K-means (Wagstaff et al., 2001) and Gaussian mixture models (Shental et al., 2003; Law et al., 2004, 2005) have been proposed to consider must-link and cannot-link relationships during clustering. A representative example is PCKMeans (Basu et al., 2004), which extends K-means by incorporating pairwise constraints into its objective function. However, these methods often struggle with robustness in high-dimensional embeddings or noisy datasets. Spectral methods represent another major branch, where constraints are encoded in graph structures (Kamvar et al., 2003; Kulis et al., 2005) or eigenspace manipulations (De Bie et al., 2004). These methods offer theoretically-grounded solutions but require a significant percentage of constraints and may overfit noise or inconsistencies in the constraint set. Among these, Constrained Spectral Clustering (CSC(Wang et al., 2014) stands out by integrating pairwise constraints into the spectral embedding through modifications to the graph Laplacian. Deep learning-based approaches address these limitations by leveraging neural networks to learn feature representations that respect constraints (Hsu and Kira, 2015; Chen, 2015). Methods such as SDEC (Ren et al., 2019) and constrained IDEC (Zhang et al., 2021c) refine latent spaces for clustering, but their discriminative nature limits their ability to model underlying data distributions. Generative extensions remain underexplored, with only a few studies (Luo et al., 2018) attempting to integrate constraints into probabilistic frameworks, often at the cost of computational efficiency. Despite progress, existing methods fail to fully exploit the structure of constraints in generative settings, leaving room for approaches that unify flexibility and interpretability in constrained clustering. In order to bridge this gap, we introduce weighting schemes based on PCKMeans and CSC and propose WCKMeans and WCSC including a pointwise mutual information weighting method, which elevates the constrained clustering from simply considering two types of constraints as equally crucial to specific weights for each constraint.
@@ 3.1. Notation and Terminology
We typically use $\{\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots,\mathcal{Y}_{K}\}$ to represent a set of $K$ disjoint clusters and a cluster assignment matrix $\mathbf{Y}\in\mathbb{R}^{n\times K}$ to symbolize the text-cluster memberships. In particular, $\mathbf{Y}_{i,k}=1$ if $t_{i}\in\mathcal{Y}_{k}$ and 00 otherwise. A must-link (resp. cannot-link) refers to a pair of text instances $t_{i},t_{j}\in\mathcal{T}$ that are required to be in the same cluster (resp. distinct clusters). We use $[1,n]$ to denote the set of integers $\{1,2,\ldots,n\}$.
@@ 3.2. Problem Statement
4. Synoptic Overview of TECL
Figure 1. An overview of TECL.
This section provides an overview of our proposed TECL framework for TCCL. As illustrated in Fig. 1, at a high level, TECL works in a cascaded pipeline, where the first step seeks to construct MLS $\mathcal{M}$ and CLS $\mathcal{C}$, i.e., must-links or cannot-links, via our text pair/triplet queries to LLMs, while the second step produces the cluster assignments for all text instances using the weighted constrained clustering algorithms that consider the importance weight of each constraint in $\mathcal{M}$ and $\mathcal{C}$.
Generating $\mathcal{M}$ and $\mathcal{C}$ via LLMs. More precisely, TECL relies on EdgeLLM or TriangleLLM to create MLS $\mathcal{M}$ and CLS $\mathcal{C}$, which build text pairs or triplets to query the LLMs to form pair-wise constraints. We refer to these two types of queries as edge and triangle queries henceforth, respectively. Denote by $s$ the average amount of tokens for a single text in $\mathcal{T}$. In EdgeLLM, we first generate
(3) $N={Q}/{(2s)}$
@@ 4. Synoptic Overview of TECL
In WCSC, we first normalize the PMI values of pairs in $\mathcal{M}$ and $\mathcal{C}$ into a fixed range (e.g., $[0.5,1.5]$) and then construct the constraint matrix $\boldsymbol{\Theta}$ where $\boldsymbol{\Theta}_{a,b}=\textsf{PMI}(t_{a},t_{b})$, $-\textsf{PMI}(t_{a},t_{b})$ or 00 when $(t_{a},t_{b})$ is a must-link, cannot-link, or has no constraints, respectively. $\boldsymbol{\Theta}$ is later used to build the graph Laplacian matrix for subsequent spectral clustering.
@@
In the succeeding Sections 5 and  6, we mainly focus on elaborating on the algorithmic details of our EdgeLLM and TriangleLLM methods for constructing MLS $\mathcal{M}$ and CLS $\mathcal{C}$.
@@ 5. The EdgeLLM Method
In this section, we first delineate our greedy algorithm in EdgeLLM for identifying the $N$ informative edges (Sections 5.1 and 5.2). Later, in Section 5.3, we introduce our carefully crafted prompts to LLMs and strategies to extract must-link and cannot-link constraints.
@@ 5.1. Spanning Edge Centrality
The SEC $\Phi(t_{a},t_{b})$ of each edge $(t_{a},t_{b})\in\mathcal{G}$ is bounded by
$\frac{1}{2}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right)\leq\Phi(t_{a},t_{b})\leq\frac{1}{1-\lambda_{2}}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right),$
where $\lambda_{2}<1$ stands for the second largest eigenvalue of the normalized adjacency matrix of $\mathcal{G}$.
In an unweighted graph $\mathcal{G}$, the SEC $\Phi(t_{a},t_{b})$ of an edge $(t_{a},t_{b})$ can be interpreted as the fraction of spanning trees of $\mathcal{G}$ that contains $(t_{a},t_{b})$. Intuitively, a large $\Phi(t_{a},t_{b})$ indicates that edge $(t_{a},t_{b})$ plays a critical role in the graph topology. However, the exact computation of $\Phi(t_{a},t_{b})$ is rather costly, and even most approximate algorithms still struggle to cope with medium-sized graphs (Spielman and Srivastava, 2008; Zhang et al., 2023b).

### v2

# Cequel: Cost-Effective Querying of Large Language Models for Text Clustering

@@ Abstract.
Text clustering aims to automatically partition a collection of documents into coherent groups based on their linguistic features. In the literature, this task is formulated either as metric clustering over pre-trained text embeddings or as graph clustering based on pairwise similarities derived from an oracle, e.g., a large machine learning model. Recent advances in large language models (LLMs) have significantly improved this field by providing high-quality contextualized embeddings and accurate semantic similarity estimates. However, leveraging LLMs at scale introduces substantial computational and financial costs due to the large number of required API queries or inference calls.
To address this issue, we propose Cequel, a cost-effective framework that achieves accurate text clustering under a limited budget of LLM queries. At its core, Cequel constructs must-link and cannot-link constraints by selectively querying LLMs on informative text pairs or triplets, identified via our proposed algorithms, EdgeLLM and TriangleLLM. These constraints are then utilized in a weighted constrained clustering algorithm to form high-quality clusters. Specifically, EdgeLLM and TriangleLLM employ carefully designed greedy selection strategies and prompting techniques to identify and extract informative constraints efficiently. Experiments on multiple benchmark datasets demonstrate that Cequel consistently outperforms existing methods in unsupervised text clustering under the same query budget.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 1787ccs: Mathematics of computing Graph algorithmsccs: Information systems Clusteringccs: Information systems Language models
@@ 1. Introduction
With the advent of LLMs, another way for text clustering is to obtain similarity scores of text pairs from LLMs to define a graph clustering problem (Silwal et al., 2023). This methodology enables us to model the correlations between text instances with highly accurate semantic similarities/dissimilarities from LLMs, and hence, producing high-quality clustering results (Silwal et al., 2023; Zhang et al., 2023a). Despite the superb performance achieved, it requires pair-wise similarities between all text instances, leading to $O(n^{2})$ inference calls to the models, where $n$ is the total number of texts in the input text corpus, which is infeasible for clustering large-scale text collections in practice. Very recently, there are a number of works (Zhang et al., 2023a; Wang et al., 2023; Yang et al., 2024; Huang and He, 2024) exploring other ways of leveraging the feedback from instruction-tuned LLMs to guide text clustering. However, all of them still demand substantial queries to LLMs (typically with a token count of over $10\times$ the corpus size) and/or rely on skeleton support from fine-tuning embedders such as BERT, Instructor, and E5, which is time-consuming.
@@
Motivated by this, this paper proposes Cequel (short for Cost-effective querying of LLMs for text clustering), which aims to maximize the text clustering quality within a fixed budget (i.e., for token consumption) for querying LLMs. To be more specific, Cequel is inspired by the constrained clustering paradigm (a.k.a. semi-supervised clustering) (Dinler and Tural, 2016), which additionally takes as input a number of constraints for pairs of data points, e.g., must-links or cannot-links, from human experts. As such, clusters can thus be obtained based on these constraints as supervision signals and cheap text embeddings from small PLMs, e.g., BERT or Sentence-BERT (Reimers, 2019). Given the human-level performance of LLMs in text comprehension, a simple and straightforward approach is to employ LLMs as the domain expert to create pair-wise pseudo constraints. However, this scheme poses two pressing challenges. First, it still remains unclear which text instances should be selected for querying and how to elicit high-caliber responses from LLMs. Secondly, the answers from LLMs might be noisy or random, engendering unreliable constraints that mislead the subsequent clustering.
@@
To tackle the above problems, Cequel includes two approaches EdgeLLM and TriangleLLM that both construct must-link and cannot-link constraints with LLMs in a cost-effective manner. EdgeLLM harnesses the prominent spanning edge centrality (Teixeira et al., 2013; Chandra et al., 1989) in graph theory to quantify the informativeness of text pairs (referred to as edges) and a well-thought-out greedy algorithm to achieve a linear-complexity identification of crucial edges. For each selected edge $(t_{a},t_{b})$, EdgeLLM elicits the relation, e.g., must-link (must be in the same cluster) or cannot-link (cannot belong to the same cluster), of $t_{a}$ and $t_{b}$ from LLMs with carefully-crafted prompts containing the task context and text contents. Based thereon, we further develop TriangleLLM that greedily selects crucial text triplets (referred to as triangles) for querying, which draws inspiration from the fact that humans represent an instance through comparing with others (Nosofsky, 2011). On top of that, Cequel refines classic constrained clustering approaches, i.e., constrained spectral clustering (Wang et al., 2014) and constrained K-Means (Wang et al., 2014), as WCSC and WCKMeans, with our carefully designed penalty weights for pairwise constraints based on the pointwise mutual information (Church and Hanks, 1990) calculated from the text embeddings, so as to eliminate the side effects of misinformation introduced by LLMs.
@@
Our extensive experiments comparing Cequel against multiple strong baselines under the same query budget for LLMs on five benchmark text datasets demonstrate the consistent and superior performance of Cequel in unsupervised text clustering tasks. In particular, on the well-known Tweet dataset, Cequel is able to attain an accuracy of $73.1\%$ with a remarkable improvement of $7.29\%$ over the state-of-the-art at a monetary cost of \$0.0185 for calling GPT-4.0-Mini.
@@ 2.1. Text Clustering
Early methods for text clustering rely on vector space models, e.g., bag-of-words and TF-IDF (Schütze et al., 2008), representing documents as high-dimensional sparse vectors. Word2Vec (Mikolov, 2013) and GloVe (Pennington et al., 2014) introduce low-dimensional dense embeddings but fail to capture contextual variations. Transformer-based language models, e.g., BERT (Devlin, 2018) and its successors (Liu, 2019; Reimers, 2019; Su et al., 2023), mitigate these issues by generating contextual embeddings, which have set new benchmarks in text clustering.
Although embedding-based methods are simple and efficient, they can fail to capture deeper semantic nuances (Sun et al., 2019; Saha, 2023; Ravi and Kulkarni, 2023; Viswanathan et al., 2023; Muennighoff et al., 2023). A series of deep clustering approaches (Zhou et al., 2022) are proposed to learn representations and cluster assignments jointly. These include multi-stage (Tao et al., 2021; Huang et al., 2014), iterative (Yang et al., 2016; Caron et al., 2018; Van Gansbeke et al., 2020; Niu et al., 2022; Chang et al., 2017; Niu et al., 2020), generative (Dilokthanakul et al., 2016), and simultaneous methods (Xie et al., 2016; Zhang et al., 2021a; Hadifar et al., 2019). While effective, these approaches typically demand training on large datasets and careful tuning.
Recently, several efforts have harnessed LLMs for enhanced text clustering. Some have validated the superiority of LLM embeddings for clustering (Keraghel et al., 2024; Petukhova et al., 2025). Others propose using LLMs to generate cluster labels or explanations (Wang et al., 2023; De Raedt et al., 2023; Huang and He, 2024). Methods like ClusterLLM (Zhang et al., 2023a) use LLM-predicted relations to guide fine-tuned embedders, while others iteratively refine clusters (Feng et al., 2024; Yang et al., 2024). The active oracle process in (Viswanathan et al., 2023) is most similar to ours. Despite encouraging results, prior works strongly rely on extensive LLM queries, costly fine-tuning, and/or post-corrections. Unlike them, our method achieves superior performance with a sufficiently low number of queries and budget, without fine-tuning or post-correction.
@@ 2.2. Constrained Clustering
To improve clustering, constrained clustering adopts side information like must-link (ML) or cannot-link (CL) constraints. Classical approaches incorporate these into traditional algorithms, such as constraint-based variants of K-means (Wagstaff et al., 2001) and Gaussian mixture models (Shental et al., 2003; Law et al., 2004, 2005). A representative example is PCKMeans (Basu et al., 2004), which extends K-means by adding pairwise constraints to its objective function. Spectral methods are another branch, encoding constraints in graph structures (Kamvar et al., 2003; Kulis et al., 2005) or eigenspace manipulations (De Bie et al., 2004). Among these, Constrained Spectral Clustering (CSC(Wang et al., 2014) integrates constraints by modifying the graph Laplacian.
Deep learning-based approaches leverage neural networks to learn feature representations that respect constraints (Hsu and Kira, 2015; Chen, 2015). Methods such as SDEC (Ren et al., 2019) and constrained IDEC (Zhang et al., 2021c) refine latent spaces for clustering. However, generative extensions remain underexplored, with few studies (Luo et al., 2018) integrating constraints into probabilistic frameworks. Despite progress, existing methods fail to fully exploit the structure of constraints in generative settings. To bridge this gap, we introduce weighting schemes based on PCKMeans and CSC, proposing WCKMeans and WCSC. This elevates constrained clustering from treating constraints as binary to assigning specific weights for each one.
@@ 3.1. Notation and Terminology
We typically use $\{\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots,\mathcal{Y}_{K}\}$ to represent a set of $K$ disjoint clusters and a cluster assignment matrix $\mathbf{Y}\in\mathbb{R}^{n\times K}$ to symbolize the text-cluster memberships. In particular, $\mathbf{Y}_{i,k}=1$ if $t_{i}\in\mathcal{Y}_{k}$ and 0 otherwise. A must-link (resp. cannot-link) refers to a pair of text instances $t_{i},t_{j}\in\mathcal{T}$ that are required to be in the same cluster (resp. distinct clusters). We use $[1,n]$ to denote the set of integers $\{1,2,\ldots,n\}$.
@@ 3.2. Problem Statement
4. Synoptic Overview of Cequel
Figure 1. An overview of Cequel.
This section provides an overview of our proposed Cequel framework for TCCL. As illustrated in Fig. 1, at a high level, Cequel works in a cascaded pipeline, where the first step seeks to construct MLS $\mathcal{M}$ and CLS $\mathcal{C}$, i.e., must-links or cannot-links, via our text pair/triplet queries to LLMs, while the second step produces the cluster assignments for all text instances using the weighted constrained clustering algorithms that consider the importance weight of each constraint in $\mathcal{M}$ and $\mathcal{C}$.
Generating $\mathcal{M}$ and $\mathcal{C}$ via LLMs. More precisely, Cequel relies on EdgeLLM or TriangleLLM to create MLS $\mathcal{M}$ and CLS $\mathcal{C}$, which build text pairs or triplets to query the LLMs to form pair-wise constraints. We refer to these two types of queries as edge and triangle queries henceforth, respectively. Denote by $s$ the average amount of tokens for a single text in $\mathcal{T}$. In EdgeLLM, we first generate
(3) N=Q/(2s)N={Q}/{(2s)}
@@ 4. Synoptic Overview of Cequel
In WCSC, we first normalize the PMI values of pairs in $\mathcal{M}$ and $\mathcal{C}$ into a fixed range (e.g., $[0.5,1.5]$) and then construct the constraint matrix $\boldsymbol{\Theta}$ where $\boldsymbol{\Theta}_{a,b}=\textsf{PMI}(t_{a},t_{b})$, $\textsf{PMI}(t_{a},t_{b})$ or 0 when $(t_{a},t_{b})$ is a must-link, cannot-link, or has no constraints, respectively. $\boldsymbol{\Theta}$ is later used to build the graph Laplacian matrix for subsequent spectral clustering.
@@
In the succeeding § 5 and  6, we mainly focus on elaborating on our EdgeLLM and TriangleLLM methods for constructing MLS $\mathcal{M}$ and CLS $\mathcal{C}$. Given space limits, some details are deferred to Appendix C.
@@ 5. The EdgeLLM Method
In this section, we first delineate our greedy algorithm in EdgeLLM for identifying the $N$ informative edges (§ 5.1 and 5.2). Later, in § 5.3, we introduce our carefully crafted prompts to LLMs and strategies to extract must-link and cannot-link constraints.
@@ 5.1. Spanning Edge Centrality
The SEC $\Phi(t_{a},t_{b})$ of each edge $(t_{a},t_{b})\in\mathcal{G}$ is bounded by $\frac{1}{2}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right)\leq\Phi(t_{a},t_{b})\leq\frac{1}{1-\lambda_{2}}\left(\frac{1}{d(a)}+\frac{1}{d(b)}\right)$, where $\lambda_{2}<1$ stands for the second largest eigenvalue of the normalized adjacency matrix of $\mathcal{G}$.
In an unweighted graph $\mathcal{G}$, the SEC $\Phi(t_{a},t_{b})$ of an edge $(t_{a},t_{b})$ can be interpreted as the fraction of spanning trees of $\mathcal{G}$ that contains $(t_{a},t_{b})$. Intuitively, a large $\Phi(t_{a},t_{b})$ indicates that edge $(t_{a},t_{b})$ plays a critical role in the graph topology. However, the exact computation of $\Phi(t_{a},t_{b})$ is rather costly, and even most approximate algorithms still struggle to cope with medium-sized graphs (Spielman and Srivastava, 2008; Zhang et al., 2023b; Yang and Tang, 2023).